[Brahms-l] Shift report 20040128 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Wed Jan 28 2004 - 23:54:58 EST
------------------------------------------------------------
Date: Wed Jan 28 23:54:58 2004, Shift: 16:00-24:00
Supervisor: R. Debbe, with: 
Log book#15 pages 44 - 44

Shift Summary:
------------------------------------------------------------
Two great stores, with a hic-up in between

2. we have finished A1219 FS at 8 degrees and we are now collecting data at B608 setting. MRS continues at 40 degrees
A2255

3.  H2 is missing two TDCs 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Jan 28 23:55:14 2004

This archive was generated by hypermail 2.1.8 : Wed Jan 28 2004 - 23:55:35 EST